# Rational motives on pro-algebraic stacks

### v3

@@ 3.1 Motivic cohomology
In the following, we fix a morphism $f\colon(x,X)\rightarrow(y,Y)$ of monoidal $-{\mathcal{I}}$-algebraic stacks over $S$. Let $h\colon(x,X)\rightarrow S$ and $g\colon(y,Y)\rightarrow S$ be the morphism induced by the structure morphisms, where we view $S$ as a constant $-{\mathcal{I}}$-algebraic stack. We further assume that either $f$ is cartesian or that $f$ is adjointable and ${\mathcal{I}}\simeq\mathbb{N}_{0}$.
We fix the notation as in Remark 2.21, i.e. we have an equivalence
we denote for any $i\in{\mathcal{I}}$ by
the canonical inclusion resp. projections and we set $\gamma_{i}^{X}\coloneqq h_{X}\circ\textup{ins}_{i}^{X}$ (analogously for $(y,Y)$).
By Theorem 5, we have adjunctions $f^{*}\dashv f_{*}$ and $f_{!}\dashv f^{!}$ between $\textup{DM}(x,X)$ and $\textup{DM}(y,Y)$ satisfying all the properties stated in the theorem. Note that even though $(x,X)$ and $(y,Y)$ are not assumed to be strict, we still have adjunctions $h^{*}\dashv h_{*}$ and $g^{*}\dashv g_{*}$ by Lemma 2.22.
@@
Let $f\colon(x,X)\rightarrow(y,Y)$ and $g\colon(y,Y)\rightarrow S$ be as above. Then we define the relative motivic cohomology of $(x,X)$ with coefficients in $M$ as
$R\Gamma(X,M)\coloneqq\underline{\operatorname{Hom}}_{{\mathcal{D}}(\mathbb{Q})}(f_{!}f^{!}1_{(x,X)},M).$
Further, we define for $n,m\in\mathbb{Z}$ the absolute motivic cohomology of $(y,Y)$ in degree $(n,m)$ by
@@
By construction of $g^{*}$ in Lemma 2.22, we have $\mathop{\textup{colim}}_{i\in{\mathcal{I}}}\textup{ins}_{i}^{Y}g^{*}_{i}$. Using adjunctions this shows that
$g_{*}\simeq\lim_{i}g_{i*}p^{Y}_{i}h_{Y}.$
@@
$g_{*}1_{(y,Y)}\simeq\lim_{i}g_{i*}p^{Y}_{i}\gamma_{0}^{Y}1_{Y_{0}}$ $\simeq\lim_{i}\mathop{\textup{colim}}_{i\rightarrow j}g_{i*}y_{ij*}y^{*}_{0j}1_{Y_{0}}$
$\simeq\lim_{i}\mathop{\textup{colim}}_{i\rightarrow j}g_{j*}1_{Y_{j}}$
$\simeq\lim_{i}\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{j*}1_{Y_{j}}$
@@
Under the notation above, Lemma 3.7 immediately shows
@@
We will see below, that if $(y,Y)$ is classical and $Y\rightarrow S$ is of finite type, then this colimit presentation of the absolute motivic cohomology of $Y$ agrees with the classical motivic cohomology of $Y$. In this setting, we can also look at the functor induced by the usual *-pushforward on Artin-stacks $\textup{DM}(y,Y)\simeq\textup{DM}(Y)\rightarrow\textup{DM}(S)$. This agrees by construction with $g_{*}\colon\textup{DM}(y,Y)\rightarrow\textup{DM}(S)$.
Further, in Example 3.12 and 3.13, we can use this computation for the absolute motivic cohomology, even in the non-finite type case (see the examples for an elaborate discussion).
Lemma 3.9.
Assume $(y,Y)$ is classical such that $g\colon Y\rightarrow S$ is locally of finite type. Let us denote by $\alpha\colon\textup{DM}(y,Y)\xrightarrow{\sim}\textup{DM}(Y)$ the equivalence induced by Proposition 3.1. Let us denote by $\tilde{g}\colon\textup{DM}(Y)\rightarrow\textup{DM}(S)$ the usual *-pushforward. Then we have
$g_{*}1_{(y,Y)}\simeq\tilde{g}\alpha 1_{(y,Y)}.$
Proof.
By construction, the last right hand side of Lemma 3.7 is equivalent to $\tilde{g}\alpha 1_{(y,Y)}$. ∎
@@
(cf. Lemma 3.7). Therefore, by continuity of DM (cf. [CD19, Thm. 14.3.1]) and descent, we have
@@
Let us remark that continuity of DM (cf. [CD19, Thm. 14.3.1]) shows that if $(y,Y)$ is a pro system of finite type $S$-schemes such that $Y$ is representable by a finite type $S$-scheme, we have again $H^{n,m}(Y,\mathbb{Q})\cong\operatorname{Hom}_{h\textup{DM}(Y)}(1_{Y},1_{Y}(n)[m])$.
Example 3.12.
[…]
where the second isomorphism follows from direct computations on the level of cycles (cf. [Bro16, Lem. 1.4.6 (i)]).
Example 3.13.
Let $Z\rightarrow S$ be an integral $S$-scheme with function field $K$. Let us consider the family $(U)_{U\in{\mathcal{U}}_{K}}$ of all affine open subschemes of $Z$. If we denote the respective inclusions by $\iota_{UU^{\prime}}\colon U^{\prime}\hookrightarrow U$, then this family assembles to a classical monoidal pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota,\operatorname{Spec}(K))$. Let $T\rightarrow Z$ be a flat morphism of $S$-schemes. Then base change of $T$ along the system $(\iota,\operatorname{Spec}(K))$ induces a classical monoidal pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota_{T},T_{K})$. Using Remark 3.6, we have
where the second isomorphism follows from direct computations on the level of cycles (cf. [Bro16, Lem. 1.4.6 (ii)]).
4 Applications
[…]
Next, we will apply our machinery to the stack of displays Disp (cf. Section 4.2). The stack Disp can be expressed as a strict monoidal $-\mathbb{N}$-algebraic stack $(\tau,\textup{Disp})$. Thus, the underling motive of $(\tau,\textup{Disp})$ is equivalent to the motive of its $1$-truncation. In particular, this allows us to compute the absolute motivic cohomology of Disp and by comparison also of the stack of Barsotti-Tate groups. Further, by earlier computation of the author, we see that the motive of Disp inside $\textup{DM}(\mathbb{F}_{p})$ is Tate.
@@ 4.1 Action of the absolute Galois group on the motivic homology spectrum
[…]
$f_{\bullet}\colon(x,{\mathfrak{X}}_{L})\rightarrow(q,L)$
@@
$\textup{DM}(x,{\mathfrak{X}}_{L})\simeq\mathop{\textup{colim}}_{{\mathcal{G}}^{L}_{K},x^{*}}\textup{DM}(x,{\mathfrak{X}}_{L}).$
Thus, to understand the action of $\operatorname{Gal}(L/k)$ on $f_{L*}1_{{\mathfrak{X}}_{L}}$, we may restrict $(x,{\mathfrak{X}}_{L})$ and $(q,L)$ to ${\mathcal{G}}^{L}_{k}$ and work with their underlying pro-${\mathcal{G}}^{L}_{k}$-algebraic stacks. By abuse of notation, we will keep the notation of $(x,{\mathfrak{X}}_{L})$ and $(q,L)$.
@@
$\textup{DM}(x,{\mathfrak{X}}_{L})\simeq\textup{DM}({\mathfrak{X}}_{L})\textup{and }\textup{DM}(q,L)\simeq\textup{DM}(L).$
We will see in the next lemma that our formalism enables us to compute the motive $f_{L*}1_{{\mathfrak{X}}_{L}}$ by "pulling back" $f_{*}1_{{\mathfrak{X}}}$ along the natural map $\textup{DM}(k)\rightarrow\textup{DM}(q,L)$.
@@
$f_{L*}1_{{\mathfrak{X}}}\simeq f_{L*}x_{L/k}^{*}1_{{\mathfrak{X}}}\simeq q_{L/k}^{*}f_{*}1_{{\mathfrak{X}}}.$
As expected this also holds by the lemma below, if $L/k$ is not finite. If $L/k$ is not finite this does not hold as there is no base change formalism that shows this result. In our formalism we have an analog of a smooth base change in this case as in Remark 3.3.
Next, we claim that $\operatorname{Gal}(L/k)$ acts on $h_{*}1_{{\mathfrak{X}}_{L}}$. To see this, it is enough to construct actions of $\operatorname{Gal}(E/k)$ on $h_{*}1_{{\mathfrak{X}}_{L}}$ for any finite Galois extension $E/k$ that is compatible with the presentation of $\operatorname{Gal}(L/k)$ as a limit of such.
@@
$\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{k}^{L}}h_{*}1_{{\mathfrak{X}}}^{\operatorname{Gal}(E/k)}\simeq\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{k}^{L}}\mathop{\textup{colim}}_{F\in{\mathcal{G}}_{k}^{L}}q_{F/k*}f_{F*}1_{{\mathfrak{X}}_{F}}^{\operatorname{Gal}(E/k)}\simeq f_{*}1_{{\mathfrak{X}}}.$
(cf. [Ayo07, Lem. 2.1.166]). In particular, for ${\mathfrak{X}}=[X/G]$ with $X$ smooth $k$-scheme and $G$ linear algebraic $k$-group, we have
$A^{n}_{G}(X,m)_{\mathbb{Q}}\simeq A^{n}_{G_{L}}(X_{L},m)^{\operatorname{Gal}(L/k)}_{\mathbb{Q}}$
by Example 3.12. This recovers the computation on the level of cycles purely formally (cf. [Bro18, Lem. 1.3.6]).

### v4

@@ 4 Motivic cohomology of pro-algebraic stacks
In the following, we fix a morphism $f\colon(x,X)\rightarrow(y,Y)$ of pro-$\mathcal{I}$}-algebraic stacks over $S$. Let $h\colon(x,X)\rightarrow S$ and $g\colon(y,Y)\rightarrow S$ be the morphism induced by the structure morphisms. We further assume that either $f$ is cartesian or that $\mathcal{I}$$\simeq\mathbb{N}$ and $f$ is adjointable.
By Theorem 2.9, we have adjunctions $f^{*}\dashv f_{*}$ and $f_{!}\dashv f^{!}$ between $\textup{DM}(x,X)$ and $\textup{DM}(y,Y)$ satisfying all the properties stated in the theorem. Note that even though $(x,X)$ and $(y,Y)$ are not assumed to be strict, we still have adjunctions $h^{*}\dashv h_{*}$ and $g^{*}\dashv g_{*}$ by Lemma 3.9.
@@
Let $f\colon(x,X)\rightarrow(y,Y)$ and $g\colon(y,Y)\rightarrow S$ be as above. Then we define the relative motivic cohomology of $(x,X)$ with coefficients in $M\in\textup{DM}(y,Y)$ as
$R\Gamma(X,M)\coloneqq\underline{\operatorname{Hom}}_{{\mathcal{D}}(\mathbb{Q})}(f_{!}f^{!}1_{(y,Y)},M)\in\textup{DGCat}_{\textup{cont}}.$
Further, we define for $n,m\in\mathbb{Z}$ the absolute motivic cohomology of $(y,Y)$ in degree $(n,m)$ as the $\mathbb{Q}$-vector space
@@
Assumption 4.3.
From now on let us assume that $\mathcal{I}$} admits an initial object $0\in{\mathcal{I}}$
@@
By construction, we have $g^{*}\simeq\textup{ins}_{0}^{Y}g^{*}_{0}$. Using adjunctions this shows that
$g_{*}\simeq g_{0*}p^{Y}_{0}h_{Y}.$
@@
$g_{*}1_{(y,Y)}\simeq g_{0*}p^{Y}_{0}\gamma_{0}^{Y}1_{Y_{0}}$ $\simeq\mathop{\textup{colim}}_{j\in{\mathcal{I}}}g_{0*}y_{0j*}y^{*}_{0j}1_{Y_{0}}$
@@
Lemma 4.4 immediately shows that for $S$ regular and $(y,Y)\rightarrow S$ smooth, we have
@@
We will see below, that if $(y,Y)$ is classical and $Y\rightarrow S$ is locally of finite type, then this colimit presentation of the absolute motivic cohomology of $Y$ agrees with the classical motivic cohomology of $Y$. In this setting, we can also look at the functor induced by the usual *-pushforward on Artin-stacks $\textup{DM}(y,Y)\simeq\textup{DM}(Y)\rightarrow\textup{DM}(S)$. This agrees by construction with $g_{*}\colon\textup{DM}(y,Y)\rightarrow\textup{DM}(S)$.
@@
(cf. Lemma 4.4). Therefore, by continuity of DM [CD19, Thm. 14.3.1] and descent, we have
@@
Example 4.7.
Let us consider the situation of Example 2.7. So let $L/K$ be an algebraic extension of fields and consider a smooth scheme $Z\rightarrow\operatorname{Spec}(K)$. We denote by ${\mathcal{F}}^{L}_{K}$ the set of finite field extensions $E/K$ contained in $L$. We denote by $(\iota,\operatorname{Spec}(L))$ the associated classical pro-${\mathcal{F}}^{L}_{K}$-algebraic stack over $K$ and by $(z,Z_{L})$ the induced classical pro-${\mathcal{F}}^{L}_{K}$-algebraic stack over $K$. Using Remark 4.2, we have
where the second isomorphism follows from direct computations on the level of cycles [Bro16, Lem. 1.4.6 (i)].
Example 4.8.
Assume that $S$ is regular. Let $Z\rightarrow S$ be a smooth integral $S$-scheme with function field $K$. Let us consider the family $(U)_{U\in{\mathcal{U}}_{K}}$ of all affine open subschemes of $Z$. If we denote the respective inclusions by $\iota_{UU^{\prime}}\colon U^{\prime}\hookrightarrow U$, then this family assembles to a classical pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota,\operatorname{Spec}(K))$. Let $T\rightarrow Z$ be a flat morphism of $S$-schemes. Then base change of $T$ along the system $(\iota,\operatorname{Spec}(K))$ induces a classical pro-${\mathcal{U}}_{K}$-algebraic stack $(\iota_{T},T_{K})$. Using Remark 4.2, we have
where the second isomorphism follows from direct computations on the level of cycles [Bro16, Lem. 1.4.6 (ii)].
4.1 Motives on Disp
Let us fix a prime $p>0$ and assume from now on that $S=\operatorname{Spec}(\mathbb{F}_{p})$. We will denote the Frobenius with $\sigma$. In the following section, we want to define the pro-algebraic stack of (truncated) Displays over $S$, after the construction of Lau (cf. [Lau13]) and show that its underlying motive in $\textup{DM}(S)$ is Tate. We do not want to go into the detail of the construction of (truncated) Displays and its relation to the stack of Barsotti-Tate groups, but rather give an equivalent definition following [Bro18].
$\begin{pmatrix}A&B\\ C&D\end{pmatrix}$
with $A\in\operatorname{GL}_{h-d}(W_{n}(R))$, $B\in\operatorname{End}(W_{n}(R)^{d},W_{n}(R)^{h-d})$, $C\in\operatorname{End}(W_{n}(R)^{h-d},I_{n+1}(R)^{d})$, where $I_{n+1}(R)$ denotes the image of the Verschiebung $W_{n+1}\rightarrow W_{n}$ and $D\in\operatorname{GL}_{d}(W_{n}(R))$. Let us denote by $G^{h,d}_{n}$ the group scheme representing this functor. Then $G^{h,d}_{n}$ acts on $X^{h}_{n}$ via $M\cdot x\coloneqq Mx\sigma^{\prime}(M)^{-1}$, where
Definition 4.9.
$\textup{Disp}_{n}^{h,d}\coloneqq\left[X^{h}_{n}/G^{h,d}_{n}\right].$
Further, we define the stack of $n$-truncated displays as
$\textup{Disp}_{n}\coloneqq\coprod_{0\leq d\leq h}\textup{Disp}_{n}^{h,d}.$
Notation 4.10.
For each $n\in\mathbb{N}$, there exists a truncation map $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$. We denote by $(\tau,\textup{Disp})$ the pro-algebraic stack induced by the $\tau_{n}$.
Lemma 4.11 ([Bro18, Lem. 2.2.2]).
@@ 4.1 Motives on Disp
Proposition 4.13.
The pro-algebraic stack $(\tau,\textup{Disp})$ is strict and tame.
Proof.
$\left[X^{h}_{n}/G^{h,d}_{n}\right]\xrightarrow{a}\left[X^{h}_{m}/G^{h,d}_{n}\right]\xrightarrow{b}\left[X^{h}_{m}/G^{h,d}_{m}\right].$
We now claim that $a^{*}$ is fully faithful (the idea is the same as in [Bro18, Thm 2.3.1]). By definition the map $X^{h}_{n}\rightarrow X^{h}_{m}$ is a $K^{h,0}_{n,m}$-torsor. By étale descent we may replace the morphism $a$ by the projection $K^{h,0}_{n,m}\rightarrow S$. By Lemma 4.11 this group is split unipotent, so it admits a normal series with successive quotients given by a vector bundle. Thus, applying again étale descent and induction, we may assume that $K^{h,0}_{n,m}$ is a vector bundle over $S$, showing that $a^{*}$ is fully faithful.
The factorization also shows that $\tau$ is tame. ∎
Lemma 4.14 ([Lau21]).
The functors $X^{h,d}_{\infty}\coloneqq\lim_{n}X^{h,d}_{n}$ and $G^{h,d}_{\infty}\coloneqq\lim_{n}G^{h,d}_{n}$ are representable by affine $\mathbb{F}_{p}$-group schemes. Moreover, $G^{h,d}_{\infty}$ acts on $X^{h,d}_{\infty}$ and we have an equivalence of algebraic stacks $\left[X^{h,d}_{\infty}/G^{h,d}_{\infty}\right]\cong\textup{Disp}$.
Proposition 4.15.
The natural morphism $\textup{DM}(\tau,\textup{Disp})\rightarrow\textup{DM}(\textup{Disp})$ induces an equivalence of DG-categories.
Proof.
Recall that $\textup{Disp}_{n}^{h,d}\cong\left[X_{n}^{h,d}/G^{h,d}_{n}\right]$. Let us consider the notation of Lemma 4.14. The pro-algebraic group scheme $G^{h,d}_{\infty}$ acts on each $X_{n}^{h,d}$ via restriction for each $n\in\mathbb{N}$.
Note that we have for any $1\leq n\leq m\leq\infty$ a cartesian diagram
(4.15.1) $X_{m}^{h,d}$$X_{n}^{h,d}$$\left[X_{m}^{h,d}/G^{h,d}_{\infty}\right]$$\left[X_{n}^{h,d}/G^{h,d}_{\infty}\right].$}
[Wed23, Cor. A.13]. In particular, the induced map
$\lim_{n}X_{n}^{h,d}\cong X_{\infty}^{h,d}\rightarrow\left[X_{\infty}^{h,d}/G^{h,d}_{\infty}\right]$
is an effective epimorphism of étale stacks and the associated Čech nerve is given by base change of the Čech nerve of $X_{1}^{h,d}\rightarrow\left[X_{1}^{h,d}/G^{h,d}_{\infty}\right]$.
Each transition map $\left[X_{n+1}^{h,d}/G^{h,d}_{\infty}\right]\rightarrow\left[X_{n}^{h,d}/G^{h,d}_{\infty}\right]$ is an affine bundle as explained in the proof of Proposition 4.13. We also note that $G^{h,d}_{\infty}$ is an affine faithfully flat group scheme.
We have
$\textup{DM}([X_{n}^{h,d}/G^{h,d}_{\infty}])\simeq\lim_{n}(\textup{DM}(X_{n}^{h,d})$$\textup{DM}(X_{n}^{h,d}\times_{\mathbb{F}_{p}}G^{h,d}_{\infty})$}),{...),}
for all $1\leq n\leq\infty$ [RS20, Thm. 2.2.16]. Having the cartesian diagram (4.15.1) in mind, we can now repeat the proof of Proposition 2.13 and obtain
$\textup{DM}(\left[X_{\infty}^{h,d}/G^{h,d}_{\infty}\right])\simeq\mathop{\textup{colim}}_{n}\textup{DM}(\left[X_{n}^{h,d}/G^{h,d}_{\infty}\right]).$
Finally, by [RS20, Prop. 2.2.11], we have
$\textup{DM}(\left[X_{n}^{h,d}/G^{h,d}_{\infty}\right])\simeq\textup{DM}(\left[X_{n}^{h,d}/G^{h,d}_{n}\right])$
concluding the proof. ∎
Theorem 4.16.
The pullback along the truncation map $\textup{Disp}\rightarrow\textup{Disp}_{1}$ induces a fully faithful embedding.
$\textup{DM}(\textup{Disp}_{1})\hookrightarrow\textup{DM}(\textup{Disp})$
In particular, we have an equivalence of motives
$M(\textup{Disp})\simeq M(\textup{Disp}_{1})$
inside $\textup{DM}(\mathbb{F}_{p})$ and moreover it is contained in the full stable cocomplete subcategory of $\textup{DM}(\mathbb{F}_{p})$ generated by Tate motives.
Proof.
By Proposition 4.13 the natural map $\textup{DM}(\textup{Disp}_{1})\rightarrow\textup{DM}(\tau,\textup{Disp})$ is fully faithful. By Proposition 4.15 the natural map $\textup{DM}(\tau,\textup{Disp})\rightarrow\textup{DM}(\textup{Disp})$ is an equivalence. This proves that $\textup{DM}(\textup{Disp}_{1})\rightarrow\textup{DM}(\textup{Disp})$ is fully faithful and $M(\textup{Disp})\simeq M(\textup{Disp}_{1})$. The fact, that this motive is Tate follows [^4] Note that $\textup{Disp}_{1}$ is isomorphic to the stack of $F$-zips, considered in [Yay23]. from [Yay23]. ∎
Remark 4.17.
Brokemper computes the Chow groups of $\textup{Disp}_{1}^{h,d}$ explicitly in [Bro18, Thm. 2.3.3]. Thus, Theorem 4.16 and Remark 2.15 tell us that the motivic cohomology of Disp can be computed by the Chow groups of $\textup{Disp}_{1}^{h,d}$. To be more precise, we can compute
Remark 4.18.
Let BT be the stack of Barsotti-Tate groups (short, BT-groups) over $S$. For any $n\in\mathbb{N}$, we denote by $\textup{BT}_{n}$ the stack of level-$n$ BT-groups. For any $n\geq 0$ there exists a truncation map $\tau_{n}\colon\textup{BT}_{n+1}\rightarrow\textup{BT}_{n}$ and $\lim_{n,\tau}\textup{BT}_{n}\simeq\textup{BT}$. The maps $\tau_{n}$ are smooth (cf. [Ill85, Thm. 4.4]). Note that every level-$n$ BT-group admits the notion of a height and dimension, which are locally constant functions over $S$. In particular, we can write $\textup{BT}_{n}\simeq\coprod_{0\leq d\leq h}\textup{BT}^{h,d}_{n}$, where $\textup{BT}^{h,d}_{n}$ denotes the substack of BT generated by BT-groups of dimension $d$ and height $h$.
We want to remark, that there is a morphism $\phi\colon\textup{BT}\rightarrow\textup{Disp}$ compatible with truncations (cf. [Lau13]). Moreover, each of the maps $\phi_{n}$ is smooth and an equivalence on geometric points (cf. [Lau13, Thm. A]). In particular, $\phi_{n}$ is a universal homeomorphism. Note however, that this is not enough to see that $\textup{DM}(\tau,\textup{Disp})\simeq\textup{DM}(\tau,\textup{BT})$ as the morphisms $\phi_{n}$ are not representable. We, conjecture that $\phi^{*}\colon\textup{DM}(\textup{Disp})\rightarrow\textup{DM}(\textup{BT})$ restricts to Tate-motives proving that the underlying motive of BT and Disp are equivalent. However, as pro-algebraic stacks the motivic cohomology groups of BT and Disp are isomorphic.
Corollary 4.19.
We have an isomorphism of motivic cohomology groups
$H^{*,*}(\textup{BT},\mathbb{Q})\cong H^{*,*}(\textup{Disp},\mathbb{Q})\cong H^{*,*}(\textup{Disp}_{1},\mathbb{Q}).$
Proof.
This follows form the definition of the motivic cohomology for the pro-algebraic stack $(\tau,\textup{BT})$, Theorem 4.16 and [Bro18, Thm. 2.5.4]. ∎
@@ 4.2 Action of the absolute Galois group on the motivic homology spectrum
Let us come back to the setting of Example 2.7 and fix the setting for this subsection. So as in the example, let $L/K$ be a Galois extension of fields. Let $X$ be a smooth $K$-scheme. We denote by ${\mathcal{F}}^{L}_{K}$ the filtered category of all finite extensions of $K$ contained in $L$. The inclusion along finite extensions $E\hookrightarrow F$ of $K$ yield pro-${\mathcal{F}}^{L}_{K}$-algebraic stacks $(x,X_{L})$ and $(q,L)$. The base change of the structure map $f\colon X\rightarrow\operatorname{Spec}(K)$ to $E\in{\mathcal{F}}^{L}_{K}$ yields a map of diagrams
$f_{\bullet}\colon(x,X_{L})\rightarrow(q,L)$
@@
$\textup{DM}(x,X_{L})\simeq\mathop{\textup{colim}}_{E\in{\mathcal{G}}^{L}_{K},x^{*}}\textup{DM}(x,X_{E}).$
Thus, to understand the action of $\operatorname{Gal}(L/K)$ on $f_{L*}1_{(x,X_{L})}$, we may restrict $(x,X_{L})$ and $(q,L)$ to ${\mathcal{G}}^{L}_{K}$ and work with their underlying pro-${\mathcal{G}}^{L}_{K}$-algebraic stacks. By abuse of notation, we will keep the notation of $(x,X_{L})$ and $(q,L)$.
@@
$\textup{DM}(x,X_{L})\simeq\textup{DM}(X_{L})\textup{ and }\textup{DM}(q,L)\simeq\textup{DM}(L)$
We will see in the next lemma that our formalism enables us to compute the motive $f_{L*}1_{(x,X_{L})}$ by "pulling back" $f_{*}1_{X}$ along the natural map $\textup{DM}(K)\rightarrow\textup{DM}(q,L)$.
@@
$f_{L*}1_{(x,X_{L})}\simeq f_{L*}x_{L/K}^{*}1_{X}\simeq q_{L/K}^{*}f_{*}1_{X}.$
If $L/K$ is not finite this does not hold as there is no base change formalism that shows this result. In our formalism we have an analog of a smooth base change in this case as in Remark 3.21.
Next, we claim that $\operatorname{Gal}(L/K)$ acts on $h_{*}1_{(x,X_{L})}$. To see this, it is enough to construct actions of $\operatorname{Gal}(E/K)$ on $h_{*}1_{(x,X_{L})}$ for any finite Galois extension $E/K$ that is compatible with the presentation of $\operatorname{Gal}(L/K)$ as a limit of such.
@@
$\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{K}^{L}}h_{*}1_{X}^{\operatorname{Gal}(E/K)}\simeq\mathop{\textup{colim}}_{E\in{\mathcal{G}}_{K}^{L}}\mathop{\textup{colim}}_{F\in{\mathcal{G}}_{E}^{L}}q_{F/K*}f_{F*}1_{X_{F}}^{\operatorname{Gal}(E/K)}\simeq f_{*}1_{X},$
where we use [Ayo07, Lem. 2.1.166] in the first equivalence. In particular, for a $X$ smooth $K$-scheme, we have
$A^{n}(X,m)_{\mathbb{Q}}\simeq A^{n}(X_{L},m)^{\operatorname{Gal}(L/K)}_{\mathbb{Q}}$
by Example 4.7. This recovers the computation on the level of cycles [Bro18, Lem. 1.3.6].